1. A golf ball of mass 0.04kg hit off the tree at a speed of 40.5 m/s. The golf club was in contact with the ball for 0.003s. Find the average force exerted on the ball by the golf club.
  2. A golf ball of mass 0.08kg hit off the tree at a speed of 38m/s. The golf club was in contact with the ball for 0.003s. Find the impulse imparted to the golf ball.
  3. A flywheel is initially rotating at 5.9 rad/s and has a constant angular acceleration. After 3s it has rotated through 30 rad. The magnitude of its angular acceleration in rad/sis:
1
Expert's answer
2022-01-07T09:28:57-0500

1. "F=mv/t=540\text{ N}."

2. "I=mv=3.04\text{ kg m/s}."

3. "30=5.9·3+\alpha·3^2/2→\alpha=2.73\text{ rad/s}^2."